OpenStudy (waheguru):

i really need help what is 4 2/3 + 3 1/2

OpenStudy (waheguru):

i am getting 8 1/6 but my cal shows 7 7/6 ???

OpenStudy (anonymous):

(4/1)+(2/3)+(3/1)+(1/2)=(24+4+18+3)/6=49/6=(48/6)+(1/6) =8 1/6

OpenStudy (anonymous):

it is just: 14/3 + 7/2 (28+21)/6 49/6 or 8 1/6

OpenStudy (anonymous):

you are correct

OpenStudy (waheguru):

if u have a calculator do the problem and u get 6 7/6?

OpenStudy (anonymous):

8 1/6 is equal to 7 7/6

OpenStudy (waheguru):

yea but what would be the right answer on a test?

OpenStudy (anonymous):

8 1/6 is right

OpenStudy (anonymous):

agree. you can still simplify 7 7/6

OpenStudy (waheguru):

ok?

OpenStudy (anonymous):

7 7/6 = 8 1/6 (which you can't break down any more. You're done.)